OpenStudy (anonymous):

Find an equation of the line that is both normal to the given parabola and also parallel to the given line. y = 4 x^2 - 7 x + 5 x - 7 y = 6

OpenStudy (anonymous):

let y=mx+c be the line and so, m=1/7.now wat d'u do , find the tangent line to the parabola whose slope is -7. u can proceed from here. understand ?

OpenStudy (anonymous):

tangent line is 8x-7. now what do i do?

OpenStudy (anonymous):

no the tangent line has to be of the form, y=-7x+c. check ur calculations. u'll get the answer as y=-7x+5

OpenStudy (anonymous):

i gav the tangent line equation...the point is (0,5) through which the normal is drawn.

OpenStudy (anonymous):

y=1/7x + 5 is the answer?

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

y = 1/7*x + c since this line is normal to the parabola, it will touch the point at which the parabola has a slope of -7 parabola's slope is 8x - 7, which is -7 when x = 0 hence line will go through (0,5) thus y = x/7 + 5